OpenStudy (anonymous):

simplify tan^2theta cos^2theta+ cos^2theta

OpenStudy (anonymous):

\[\tan^2\theta \cos^2\theta+\cos^2\theta\] \[(\sin^2\theta/\cos^2\theta)*\cos^2\theta+\cos^2\theta\] \[\sin^2\theta+\cos^2\theta = 1\]